Some sodium cyanide has been placed in a glass of orange juice to poison a secret agent the react...

some sodium cyanide has been placed in a glass of orange juice to poison a secret agent the reaction is

CN-(aq) + H6H7O7(aq) -> HCN(g) + C6H7O7-(aq)

i) is this reaction an acid-base reaction?
ii) would this reaction generate enough HCN to poison the agent? include the calculation of the equilibrium constant of the reaction.

CN- Kb= 2.5e-5

HC6H7O Ka= 8.4e-4
Answer #1

i)

HC6H7O7 is acid as it donates H+ and forming C6H7O-

CN- is base as it is accepting H+ and forming HCN

So, the reaction given is acid-base reaction

Answer: yes

ii)

The reaction is:

CN-(aq) + HC6H7O7(aq) -> HCN(g) + C6H7O7-(aq)

Kc = [HCN][C6H7O7-] / [CN-][HC6H7O7]

= ([HCN][OH-]/[CN-]) * [C6H7O7-][H+]/[HC6H7O7] * 1/[H+][OH-]

= ([HCN][OH-]/[CN-][H+]) * ([C6H7O7-][H+]/[HC6H7O7]) * (1/[H+][OH-])

= Kb * Ka * 1/Kw

= (2.5*10^-5)*(8.4*10^-4)*1/(1.0*10^-14)

= 2.1*10^6

Answer: 2.1*10^6